OpenStudy (shelby1290):

Do these systems have one solution, coincidental lines or infinite many solutions and why: x+y=2 3x=6-3y

OpenStudy (shelby1290):

@Directrix

Directrix (directrix):

Yikes! I meant to write divide each term by 3. Sorry.

OpenStudy (shelby1290):

3x divided by 3=1x 6 divided by 3=2 3y divided 3=1y

Directrix (directrix):

x+y=2 x = 2 + y -------- Solve the system by substitution In place of the x in x + y = 2, put (2 + y)

Directrix (directrix):

(2 + y) +y=2 and solve for y.

OpenStudy (shelby1290):

do i divide both sides by 2?

Directrix (directrix):

You can subtract 2 from both sides.

Directrix (directrix):

What do you get?

OpenStudy (shelby1290):

so 2-2? or 3-2 (2+1y=3)

Directrix (directrix):

(2 + y) +y=2 and solve for y. -2 -2 --------------- y = y

OpenStudy (shelby1290):

-4?

Directrix (directrix):

You can also show that both lines have the same slope and the same y-intercept. x+y=2 so y = -x + 2-----> and 3x=6-3y x = 2 - y so y = -x + 2 ---> The two lines are the same line so they have all points on the lines in common.

OpenStudy (shelby1290):

ohhh

Directrix (directrix):

http://www.mathwords.com/c/coincident.htm

Directrix (directrix):

Here's what I don't get about this question: Do these systems have one solution, coincidental lines or infinite many solutions One solution is out. But, the lines do coincide because of the same slope and the same y-intercept But, what's up with this: infinite many solutions ?

OpenStudy (shelby1290):

I'm going to be quoting my worksheet word for word on this one: If two lines have the same slope and the same y-intercept, the lines are called coincidental lines, and are basically the same line. If two lines are coincidental, they have infinitely many solutions

Directrix (directrix):

Tell me what you are thinking about this problem. @shelby1290

OpenStudy (shelby1290):

I just get confused when everything is out of order or when there's different steps to everything

Directrix (directrix):

The two equations are the same but one is in disguise. The first equation, each term, was multiplied by 3 to get the second equation x+y=2 3x=6-3y

Directrix (directrix):

They are two names for the same line.

Directrix (directrix):

That is why they are called coincident, because they coincide.

Directrix (directrix):

They have infinitely many solutions because there are infinitely many points on a line.

Directrix (directrix):

>>I just get confused when everything is out of order or when there's different steps to everything That's the same with all of us, really.

OpenStudy (shelby1290):

can you also solve this by putting it into y=mx+b form?

OpenStudy (shelby1290):

so then you can find out the m(slope) and the b(y-int) and figure out whether or not its one solution...coincidental etc

Directrix (directrix):

Yes. You can also show that both lines have the same slope and the same y-intercept . x+y=2 so y = -x + 2-----> and 3x=6-3y x = 2 - y so y = -x + 2 --->

OpenStudy (shelby1290):

That's how I learned to do it from people in class

Directrix (directrix):

That is a correct technique.

OpenStudy (shelby1290):

Okay there's this other question that seems a little tricky for me but I'm gonna try it out. 2y-x-4=0 3x-6y-12=0 y=mx+b 2y=-1x-4

OpenStudy (shelby1290):

does the -6y change to a positive or stay as a negative?

OpenStudy (shelby1290):

2y=-1x-4 uhhh i think the next step is to divide everything by 2 2y divided 2 is 1 -1x divided by 2 is -0.5 -4 divided by 2 is -2

Directrix (directrix):

What are the instructions for this problem?

OpenStudy (shelby1290):

they're the same as the ones that you've previously helped me with

OpenStudy (shelby1290):

I just want to know if i'm on the right track

Directrix (directrix):

2y -x -4 = 0 +x +x ------------------- 2y - 4 = x +4 +4 ------------------------- 2y = x + 4 Divide by 2 y = (1/2) x + 2 ---> list equation in y = mx + b form What did you get?

Directrix (directrix):

@shelby1290

OpenStudy (shelby1290):

am i finding y?

Directrix (directrix):

Yes, y = mx + b is what you said you wanted to do. The second equation of the second problem 3x-6y-12=0

OpenStudy (shelby1290):

-6y=3x-12

Directrix (directrix):

No

OpenStudy (shelby1290):

what is something supposed to be positive instead of negative?

Directrix (directrix):

3x-6y-12=0 -6y = -3x + 12

OpenStudy (shelby1290):

okay

Directrix (directrix):

You are skipping steps and messing up the signs.

OpenStudy (shelby1290):

what are the rules again

Directrix (directrix):

3x -6y - 12 = 0 -3x -3x ------------------------ What is that equal

OpenStudy (shelby1290):

@Directrix I'm honestly not sure. I feel like I'm wasting your time.

OpenStudy (shelby1290):

Math is my all-time worst subject

OpenStudy (shelby1290):

im so tired

Directrix (directrix):

-6y - 12 = -3x +12 +12 --------------- -6y = -3x + 12 Divide by -6 ------------------ y = (1/2) x -2 --> second equation y = (1/2) x + 2 ---> first equation same slope, different y intercept so what is the conclusion?

OpenStudy (shelby1290):

coincidental

Directrix (directrix):

No

OpenStudy (shelby1290):

my bad

Directrix (directrix):

The slopes are equal so the lines are parallel They can't be the same line because the y-intercepts are different. So, no solution.

Directrix (directrix):

>>i thought you said same slope and different y-int I did.

OpenStudy (shelby1290):

LOL sorry

Directrix (directrix):

That is how we know the lines are parallel and different lines.

Directrix (directrix):

That's a wrap, right?

OpenStudy (shelby1290):

Well, thanks a million its like 1 am where I live and I'm so exhausted. Yet, you still stayed to help me. (I already know that I'm extremely senseless in Math so it's a challenge)

OpenStudy (shelby1290):

Have a good night

Directrix (directrix):

It is 1 a.m. where I live.

Directrix (directrix):

You, too.